OpenStudy (anonymous):

Find the area of the circle. Use 3.14 Round to the nearest unit

OpenStudy (anonymous):

hold on ill tell u what I got

OpenStudy (unklerhaukus):

The area of a circle is pi times the radius squared the radius is half the diameter

OpenStudy (unklerhaukus):

What'd you get @cutegeek?

OpenStudy (anonymous):

I got 57cm2 but im unsure of that @UnkleRhaukus

OpenStudy (unklerhaukus):

should be more than that

OpenStudy (anonymous):

thought so the answer list shows 1,017 cm2 254 cm2 57 cm2 28 cm2

OpenStudy (unklerhaukus):

A = π * r^2 = π * (d/2)^2

OpenStudy (unklerhaukus):

the diameter is shown on the diagram

OpenStudy (anonymous):

I see you said more than 57 so28 is out and leaves me with 254 and 1,017cm